# LLMs Can’t Handle Peer Pressure: Crumbling under Multi-Agent Social Interactions

### v2

@@ C.2 Effect on different categories of datasets
Qwen2.5-3B Base 57.36% 59.16% +1.80% 70.27% 64.56% -5.71%
Empowered 58.86% 59.16% +0.30% 70.87% 63.66% -7.21%
Reflected 57.36% 58.26% +0.90% 70.27% 56.46% -13.81%
SFT 52.85% 60.36% +7.51% 70.27% 60.36% -9.91%
GRPO-MAS-DS-DR 62.16% 58.26% -3.90% 68.47% 64.56% -3.90%
GRPO-MAS-NS-OR 65.17% 63.96% -1.20% 71.47% 62.46% -9.01%
GRPO-nonMAS-DS-DR 60.96% 54.65% -6.31% 72.97% 63.06% -9.91%
GRPO-nonMAS-NS-OR 63.36% 60.36% -3.00% 72.37% 61.56% -10.81%
GRPO-MAS-DS-DR-LConf 61.56% 58.86% -2.70% 70.87% 60.96% -9.91%
GRPO-MAS-DS-DR-LCorr 61.56% 58.86% -2.70% 68.77% 45.65% -23.12%
GRPO-MAS-NS-OR-LConf 61.86% 61.86% +0.00% 72.97% 56.16% -16.82%
GRPO-MAS-NS-OR-LCorr 64.26% 55.56% -8.71% 72.37% 43.54% -28.83%
Qwen2.5-7B Base 68.17% 56.76% -11.41% 74.17% 68.77% -5.41%
Empowered 65.77% 57.36% -8.41% 75.08% 69.97% -5.11%
Reflected 68.17% 61.56% -6.61% 74.17% 72.97% -1.20%
SFT 60.96% 44.14% -16.82% 68.17% 50.15% -18.02%
GRPO-MAS-DS-DR 69.07% 67.87% -1.20% 71.47% 68.47% -3.00%
GRPO-MAS-NS-OR 75.08% 59.16% -15.92% 74.17% 66.07% -8.11%
GRPO-nonMAS-DS-DR 68.47% 65.17% -3.30% 71.17% 61.56% -9.61%
GRPO-nonMAS-NS-OR 76.28% 57.66% -18.62% 75.98% 54.05% -21.92%
GRPO-MAS-DS-DR-LConf 69.97% 55.56% -14.41% 72.97% 56.16% -16.82%
GRPO-MAS-DS-DR-LCorr 71.47% 57.36% -14.11% 73.27% 45.65% -27.63%
GRPO-MAS-NS-OR-LConf 73.87% 63.96% -9.91% 76.28% 59.76% -16.52%
GRPO-MAS-NS-OR-LCorr 76.88% 49.55% -27.33% 74.77% 42.04% -32.73%
Qwen2.5-14B Base 69.37% 64.56% -4.80% 74.47% 66.97% -7.51%
Empowered 70.57% 69.37% -1.20% 73.57% 69.07% -4.50%
Reflected 69.37% 64.86% -4.50% 74.47% 69.07% -5.41%
SFT 68.47% 52.85% -15.62% 70.27% 48.05% -22.22%
GRPO-MAS-DS-DR 83.18% 77.18% -6.01% 75.08% 71.77% -3.30%
GRPO-MAS-NS-OR 81.38% 78.08% -3.30% 73.27% 72.07% -1.20%
GRPO-nonMAS-DS-DR 80.48% 63.96% -16.52% 75.68% 65.77% -9.91%
GRPO-nonMAS-NS-OR 82.28% 63.96% -18.32% 76.58% 64.56% -12.01%
GRPO-MAS-DS-DR-LConf 76.88% 68.77% -8.11% 75.98% 61.56% -14.41%
GRPO-MAS-DS-DR-LCorr 76.28% 48.35% -27.93% 72.67% 42.04% -30.63%
GRPO-MAS-NS-OR-LConf 81.08% 62.46% -18.62% 76.28% 56.16% -20.12%
GRPO-MAS-NS-OR-LCorr 82.58% 48.35% -34.23% 73.57% 44.74% -28.83%
Llama3.2-3B Base 57.66% 50.75% -6.91% 63.96% 50.75% -13.21%
Empowered 59.76% 51.65% -8.11% 64.86% 55.86% -9.01%
Reflected 57.66% 49.55% -8.11% 63.96% 47.15% -16.82%
SFT 52.55% 48.65% -3.90% 57.66% 38.74% -18.92%
GRPO-MAS-DS-DR 59.16% 49.85% -9.31% 67.87% 55.26% -12.61%
GRPO-MAS-NS-OR 62.16% 60.36% -1.80% 72.07% 60.96% -11.11%
GRPO-nonMAS-DS-DR 59.46% 49.85% -9.61% 67.87% 56.46% -11.41%
GRPO-nonMAS-NS-OR 57.96% 57.36% -0.60% 73.27% 56.76% -16.52%
GRPO-MAS-DS-DR-LConf 57.66% 51.95% -5.71% 67.27% 52.25% -15.02%
GRPO-MAS-DS-DR-LCorr 57.06% 50.75% -6.31% 67.87% 45.35% -22.52%
GRPO-MAS-NS-OR-LConf 65.47% 56.46% -9.01% 71.47% 53.45% -18.02%
GRPO-MAS-NS-OR-LCorr 65.47% 56.46% -9.01% 71.47% 53.45% -18.02%
Llama3.1-8B Base 68.17% 60.66% -7.51% 73.27% 62.76% -10.51%
Empowered 65.47% 59.76% -5.71% 72.07% 65.47% -6.61%
Reflected 68.17% 50.15% -18.02% 73.27% 54.95% -18.32%
SFT 53.15% 53.45% +0.30% 65.77% 43.54% -22.22%
GRPO-MAS-DS-DR 63.66% 61.26% -2.40% 72.07% 62.76% -9.31%
GRPO-MAS-NS-OR 66.37% 63.96% -2.40% 72.37% 66.37% -6.01%
GRPO-nonMAS-DS-DR 67.27% 55.86% -11.41% 75.08% 56.76% -18.32%
GRPO-nonMAS-NS-OR 69.67% 66.07% -3.60% 73.57% 57.36% -16.22%
GRPO-MAS-DS-DR-LConf 65.77% 64.56% -1.20% 71.77% 57.36% -14.41%
GRPO-MAS-DS-DR-LCorr 65.17% 59.16% -6.01% 70.87% 46.85% -24.02%
GRPO-MAS-NS-OR-LConf 67.87% 63.96% -3.90% 70.87% 60.66% -10.21%
GRPO-MAS-NS-OR-LCorr 72.37% 63.96% -8.41% 72.07% 58.56% -13.51%
Table 9: Overall results for the Common Sense category under Kairos. Bold numbers mark per-dataset extreme (max/min) O–K $\Delta$.
@@
Qwen2.5-3B Base 37.54% 30.63% -6.91% 34.23% 44.44% +10.21%
Empowered 41.74% 30.63% -11.11% 41.14% 44.44% +3.30%
Reflected 37.54% 23.42% -14.11% 34.23% 57.06% +22.82%
SFT 54.35% 43.24% -11.11% 51.05% 72.37% +21.32%
GRPO-MAS-DS-DR 35.14% 33.63% -1.50% 44.14% 48.35% +4.20%
GRPO-MAS-NS-OR 39.04% 41.44% +2.40% 62.46% 72.07% +9.61%
GRPO-nonMAS-DS-DR 36.94% 35.44% -1.50% 46.55% 55.56% +9.01%
GRPO-nonMAS-NS-OR 44.44% 34.23% -10.21% 69.67% 63.96% -5.71%
GRPO-MAS-DS-DR-LConf 28.53% 28.53% +0.00% 40.84% 47.75% +6.91%
GRPO-MAS-DS-DR-LCorr 33.03% 33.93% +0.90% 54.05% 51.95% -2.10%
GRPO-MAS-NS-OR-LConf 38.74% 38.14% -0.60% 63.96% 60.66% -3.30%
GRPO-MAS-NS-OR-LCorr 35.44% 31.53% -3.90% 59.16% 57.06% -2.10%
Qwen2.5-7B Base 42.94% 47.45% +4.50% 74.77% 45.95% -28.83%
Empowered 44.14% 46.25% +2.10% 71.77% 45.35% -26.43%
Reflected 42.94% 42.94% +0.00% 74.77% 41.74% -33.03%
SFT 64.86% 48.65% -16.22% 60.06% 45.95% -14.11%
GRPO-MAS-DS-DR 50.45% 54.35% +3.90% 76.58% 59.46% -17.12%
GRPO-MAS-NS-OR 54.65% 55.86% +1.20% 74.77% 67.87% -6.91%
GRPO-nonMAS-DS-DR 51.05% 47.15% -3.90% 63.96% 61.56% -2.40%
GRPO-nonMAS-NS-OR 48.95% 41.14% -7.81% 73.57% 52.55% -21.02%
GRPO-MAS-DS-DR-LConf 46.55% 43.24% -3.30% 72.07% 49.55% -22.52%
GRPO-MAS-DS-DR-LCorr 43.24% 42.04% -1.20% 68.77% 54.95% -13.81%
GRPO-MAS-NS-OR-LConf 46.85% 38.74% -8.11% 74.47% 49.55% -24.92%
GRPO-MAS-NS-OR-LCorr 48.05% 41.14% -6.91% 79.28% 58.86% -20.42%
Qwen2.5-14B Base 52.25% 48.05% -4.20% 73.87% 48.95% -24.92%
Empowered 46.25% 44.74% -1.50% 68.77% 44.14% -24.62%
Reflected 52.25% 44.74% -7.51% 73.87% 52.55% -21.32%
SFT 70.87% 60.96% -9.91% 67.87% 59.76% -8.11%
GRPO-MAS-DS-DR 61.56% 59.76% -1.80% 76.58% 60.36% -16.22%
GRPO-MAS-NS-OR 60.36% 63.36% +3.00% 71.77% 61.26% -10.51%
GRPO-nonMAS-DS-DR 57.96% 53.45% -4.50% 72.07% 43.24% -28.83%
GRPO-nonMAS-NS-OR 60.66% 54.35% -6.31% 74.17% 55.86% -18.32%
GRPO-MAS-DS-DR-LConf 56.16% 52.55% -3.60% 71.47% 48.65% -22.82%
GRPO-MAS-DS-DR-LCorr 45.05% 35.74% -9.31% 78.68% 42.34% -36.34%
GRPO-MAS-NS-OR-LConf 59.16% 51.95% -7.21% 65.47% 58.26% -7.21%
GRPO-MAS-NS-OR-LCorr 58.56% 45.05% -13.51% 67.57% 47.75% -19.82%
Llama3.2-3B Base 43.54% 42.34% -1.20% 63.36% 56.76% -6.61%
Empowered 37.84% 43.54% +5.71% 69.67% 57.36% -12.31%
Reflected 43.54% 34.83% -8.71% 63.36% 56.76% -6.61%
SFT 48.05% 37.54% -10.51% 54.65% 48.95% -5.71%
GRPO-MAS-DS-DR 49.25% 43.84% -5.41% 69.97% 63.36% -6.61%
GRPO-MAS-NS-OR 51.95% 45.35% -6.61% 77.78% 61.56% -16.22%
GRPO-nonMAS-DS-DR 49.85% 39.94% -9.91% 77.18% 60.96% -16.22%
GRPO-nonMAS-NS-OR 49.55% 43.54% -6.01% 77.78% 66.37% -11.41%
GRPO-MAS-DS-DR-LConf 44.14% 40.84% -3.30% 75.08% 59.46% -15.62%
GRPO-MAS-DS-DR-LCorr 46.85% 42.34% -4.50% 77.48% 67.27% -10.21%
GRPO-MAS-NS-OR-LConf 48.05% 36.04% -12.01% 77.48% 62.16% -15.32%
GRPO-MAS-NS-OR-LCorr 48.05% 36.04% -12.01% 77.48% 62.16% -15.32%
Llama3.1-8B Base 52.55% 45.35% -7.21% 75.68% 71.47% -4.20%
Empowered 47.75% 43.54% -4.20% 76.88% 72.67% -4.20%
Reflected 52.55% 29.73% -22.82% 75.68% 33.03% -42.64%
SFT 63.36% 50.75% -12.61% 51.05% 50.15% -0.90%
GRPO-MAS-DS-DR 57.96% 51.35% -6.61% 76.88% 76.58% -0.30%
GRPO-MAS-NS-OR 59.46% 59.16% -0.30% 78.38% 72.37% -6.01%
GRPO-nonMAS-DS-DR 52.25% 45.05% -7.21% 75.98% 52.55% -23.42%
GRPO-nonMAS-NS-OR 62.16% 55.56% -6.61% 78.08% 71.17% -6.91%
GRPO-MAS-DS-DR-LConf 55.26% 48.65% -6.61% 76.58% 66.37% -10.21%
GRPO-MAS-DS-DR-LCorr 53.45% 45.65% -7.81% 77.18% 56.46% -20.72%
GRPO-MAS-NS-OR-LConf 54.65% 47.75% -6.91% 77.48% 76.88% -0.60%
GRPO-MAS-NS-OR-LCorr 61.56% 53.15% -8.41% 78.38% 66.07% -12.31%

### v3

@@ Appendix G Comprehensive Evaluation and Analysis Results
Qwen2.5-3B Base 57.36% 59.16% +3.1% 70.27% 64.56% -8.1%
Empowered 58.86% 59.16% +0.5% 70.87% 63.66% -10.2%
Reflected 57.36% 58.26% +1.6% 70.27% 56.46% -19.7%
SFT 52.85% 60.36% $\mathbf{+14.2\%}$ 70.27% 60.36% -14.1%
GRPO-MAS-DS-DR 62.16% 58.26% -6.3% 68.47% 64.56% -5.7%
GRPO-MAS-NS-OR 65.17% 63.96% -1.8% 71.47% 62.46% -12.6%
GRPO-nonMAS-DS-DR 60.96% 54.65% -10.4% 72.97% 63.06% -13.6%
GRPO-nonMAS-NS-OR 63.36% 60.36% -4.7% 72.37% 61.56% -14.9%
GRPO-MAS-DS-DR-LConf 61.56% 58.86% -4.4% 70.87% 60.96% -14.0%
GRPO-MAS-DS-DR-LCorr 61.56% 58.86% -4.4% 68.77% 45.65% -33.6%
GRPO-MAS-NS-OR-LConf 61.86% 61.86% +0.0% 72.97% 56.16% -23.1%
GRPO-MAS-NS-OR-LCorr 64.26% 55.56% -13.6% 72.37% 43.54% -39.8%
Qwen2.5-7B Base 68.17% 56.76% -16.7% 74.17% 68.77% -7.3%
Empowered 65.77% 57.36% -12.8% 75.08% 69.97% -6.8%
Reflected 68.17% 61.56% -9.7% 74.17% 72.97% -1.6%
SFT 60.96% 44.14% -27.6% 68.17% 50.15% -26.4%
GRPO-MAS-DS-DR 69.07% 67.87% -1.7% 71.47% 68.47% -4.2%
GRPO-MAS-NS-OR 75.08% 59.16% -21.2% 74.17% 66.07% -10.9%
GRPO-nonMAS-DS-DR 68.47% 65.17% -4.8% 71.17% 61.56% -13.5%
GRPO-nonMAS-NS-OR 76.28% 57.66% -24.4% 75.98% 54.05% -28.8%
GRPO-MAS-DS-DR-LConf 69.97% 55.56% -20.6% 72.97% 56.16% -23.1%
GRPO-MAS-DS-DR-LCorr 71.47% 57.36% -19.7% 73.27% 45.65% -37.7%
GRPO-MAS-NS-OR-LConf 73.87% 63.96% -13.4% 76.28% 59.76% -21.7%
GRPO-MAS-NS-OR-LCorr 76.88% 49.55% -35.5% 74.77% 42.04% $\mathbf{-43.8\%}$
Qwen2.5-14B Base 69.37% 64.56% -6.9% 74.47% 66.97% -10.1%
Empowered 70.57% 69.37% -1.7% 73.57% 69.07% -6.1%
Reflected 69.37% 64.86% -6.5% 74.47% 69.07% -7.3%
SFT 68.47% 52.85% -22.8% 70.27% 48.05% -31.6%
GRPO-MAS-DS-DR 83.18% 77.18% -7.2% 75.08% 71.77% -4.4%
GRPO-MAS-NS-OR 81.38% 78.08% -4.1% 73.27% 72.07% -1.6%
GRPO-nonMAS-DS-DR 80.48% 63.96% -20.5% 75.68% 65.77% -13.1%
GRPO-nonMAS-NS-OR 82.28% 63.96% -22.3% 76.58% 64.56% -15.7%
GRPO-MAS-DS-DR-LConf 76.88% 68.77% -10.5% 75.98% 61.56% -19.0%
GRPO-MAS-DS-DR-LCorr 76.28% 48.35% -36.6% 72.67% 42.04% -42.1%
GRPO-MAS-NS-OR-LConf 81.08% 62.46% -23.0% 76.28% 56.16% -26.4%
GRPO-MAS-NS-OR-LCorr 82.58% 48.35% $\mathbf{-41.5\%}$ 73.57% 44.74% -39.2%
Llama3.2-3B Base 57.66% 50.75% -12.0% 63.96% 50.75% -20.7%
Empowered 59.76% 51.65% -13.6% 64.86% 55.86% -13.9%
Reflected 57.66% 49.55% -14.1% 63.96% 47.15% -26.3%
SFT 52.55% 48.65% -7.4% 57.66% 38.74% -32.8%
GRPO-MAS-DS-DR 59.16% 49.85% -15.7% 67.87% 55.26% -18.6%
GRPO-MAS-NS-OR 62.16% 60.36% -2.9% 72.07% 60.96% -15.4%
GRPO-nonMAS-DS-DR 59.46% 49.85% -16.2% 67.87% 56.46% -16.8%
GRPO-nonMAS-NS-OR 57.96% 57.36% -1.0% 73.27% 56.76% -22.5%
GRPO-MAS-DS-DR-LConf 57.66% 51.95% -9.9% 67.27% 52.25% -22.3%
GRPO-MAS-DS-DR-LCorr 57.06% 50.75% -11.1% 67.87% 45.35% -33.2%
GRPO-MAS-NS-OR-LConf 65.47% 56.46% -13.8% 71.47% 53.45% -25.2%
GRPO-MAS-NS-OR-LCorr 65.47% 56.46% -13.8% 71.47% 53.45% -25.2%
Llama3.1-8B Base 68.17% 60.66% -11.0% 73.27% 62.76% -14.3%
Empowered 65.47% 59.76% -8.7% 72.07% 65.47% -9.2%
Reflected 68.17% 50.15% -26.4% 73.27% 54.95% -25.0%
SFT 53.15% 53.45% +0.6% 65.77% 43.54% -33.8%
GRPO-MAS-DS-DR 63.66% 61.26% -3.8% 72.07% 62.76% -12.9%
GRPO-MAS-NS-OR 66.37% 63.96% -3.6% 72.37% 66.37% -8.3%
GRPO-nonMAS-DS-DR 67.27% 55.86% -17.0% 75.08% 56.76% -24.4%
GRPO-nonMAS-NS-OR 69.67% 66.07% -5.2% 73.57% 57.36% -22.0%
GRPO-MAS-DS-DR-LConf 65.77% 64.56% -1.8% 71.77% 57.36% -20.1%
GRPO-MAS-DS-DR-LCorr 65.17% 59.16% -9.2% 70.87% 46.85% -33.9%
GRPO-MAS-NS-OR-LConf 67.87% 63.96% -5.7% 70.87% 60.66% -14.4%
GRPO-MAS-NS-OR-LCorr 72.37% 63.96% -11.6% 72.07% 58.56% -18.7%
Llama3.3-70B Base 78.68% 78.38% -0.4% 79.58% 75.08% -5.7%
Empowered 80.18% 80.18% +0.0% 80.48% 79.58% -1.1%
Reflected 78.68% 76.58% -2.7% 79.58% 78.98% -0.8%
Qwen2.5-32B Base 74.77% 72.07% -3.6% 79.88% 78.08% -2.3%
Empowered 74.77% 74.17% -0.8% 78.98% 78.38% -0.8%
Reflected 74.77% 74.47% -0.4% 79.88% 76.88% -3.8%
Qwen2.5-72B Base 79.88% 79.28% -0.8% 78.08% 72.97% -6.5%
Empowered 79.88% 79.58% -0.4% 78.68% 76.28% -3.1%
Reflected 79.88% 74.47% -6.8% 78.08% 77.48% -0.8%
GPT-OSS-120B Base 85.59% 72.07% -15.8% 79.88% 66.97% -16.2%
Empowered 84.68% 80.48% -5.0% 79.58% 72.37% -9.1%
Reflected 85.59% 83.18% -2.8% 79.88% 75.68% -5.3%
Gemini-2.5-Pro Base 84.38% 68.47% -18.9% 79.58% 65.77% -17.4%
Empowered 78.98% 75.38% -4.6% 78.98% 79.58% $\mathbf{+0.8\%}$
Reflected 84.38% 73.87% -12.5% 79.58% 76.28% -4.1%
GPT-5 Base 87.99% 79.88% -9.2% 80.18% 78.08% -2.6%
Empowered 86.19% 81.08% -5.9% 80.18% 78.98% -1.5%
Reflected 87.99% 81.08% -7.9% 80.18% 78.38% -2.2%
Table 10: Overall results for the Social category under Kairos. Bold numbers mark per-dataset extreme (max/min) O–K $\Delta$.
@@
Qwen2.5-3B Base 37.54% 30.63% -18.4% 34.23% 44.44% +29.8%
Empowered 41.74% 30.63% -26.6% 41.14% 44.44% +8.0%
Reflected 37.54% 23.42% -37.6% 34.23% 57.06% $\mathbf{+66.7\%}$
SFT 54.35% 43.24% -20.4% 51.05% 72.37% +41.8%
GRPO-MAS-DS-DR 35.14% 33.63% -4.3% 44.14% 48.35% +9.5%
GRPO-MAS-NS-OR 39.04% 41.44% +6.1% 62.46% 72.07% +15.4%
GRPO-nonMAS-DS-DR 36.94% 35.44% -4.1% 46.55% 55.56% +19.4%
GRPO-nonMAS-NS-OR 44.44% 34.23% -23.0% 69.67% 63.96% -8.2%
GRPO-MAS-DS-DR-LConf 28.53% 28.53% +0.0% 40.84% 47.75% +16.9%
GRPO-MAS-DS-DR-LCorr 33.03% 33.93% +2.7% 54.05% 51.95% -3.9%
GRPO-MAS-NS-OR-LConf 38.74% 38.14% -1.5% 63.96% 60.66% -5.2%
GRPO-MAS-NS-OR-LCorr 35.44% 31.53% -11.0% 59.16% 57.06% -3.5%
Qwen2.5-7B Base 42.94% 47.45% +10.5% 74.77% 45.95% -38.6%
Empowered 44.14% 46.25% +4.8% 71.77% 45.35% -36.8%
Reflected 42.94% 42.94% +0.0% 74.77% 41.74% -44.2%
SFT 64.86% 48.65% -25.0% 60.06% 45.95% -23.5%
GRPO-MAS-DS-DR 50.45% 54.35% +7.7% 76.58% 59.46% -22.4%
GRPO-MAS-NS-OR 54.65% 55.86% +2.2% 74.77% 67.87% -9.2%
GRPO-nonMAS-DS-DR 51.05% 47.15% -7.6% 63.96% 61.56% -3.8%
GRPO-nonMAS-NS-OR 48.95% 41.14% -16.0% 73.57% 52.55% -28.6%
GRPO-MAS-DS-DR-LConf 46.55% 43.24% -7.1% 72.07% 49.55% -31.2%
GRPO-MAS-DS-DR-LCorr 43.24% 42.04% -2.8% 68.77% 54.95% -20.1%
GRPO-MAS-NS-OR-LConf 46.85% 38.74% -17.3% 74.47% 49.55% -33.5%
GRPO-MAS-NS-OR-LCorr 48.05% 41.14% -14.4% 79.28% 58.86% -25.8%
Qwen2.5-14B Base 52.25% 48.05% -8.0% 73.87% 48.95% -33.7%
Empowered 46.25% 44.74% -3.2% 68.77% 44.14% -35.8%
Reflected 52.25% 44.74% -14.4% 73.87% 52.55% -28.9%
SFT 70.87% 60.96% -14.0% 67.87% 59.76% -11.9%
GRPO-MAS-DS-DR 61.56% 59.76% -2.9% 76.58% 60.36% -21.2%
GRPO-MAS-NS-OR 60.36% 63.36% +5.0% 71.77% 61.26% -14.6%
GRPO-nonMAS-DS-DR 57.96% 53.45% -7.8% 72.07% 43.24% -40.0%
GRPO-nonMAS-NS-OR 60.66% 54.35% -10.4% 74.17% 55.86% -24.7%
GRPO-MAS-DS-DR-LConf 56.16% 52.55% -6.4% 71.47% 48.65% -31.9%
GRPO-MAS-DS-DR-LCorr 45.05% 35.74% -20.7% 78.68% 42.34% -46.2%
GRPO-MAS-NS-OR-LConf 59.16% 51.95% -12.2% 65.47% 58.26% -11.0%
GRPO-MAS-NS-OR-LCorr 58.56% 45.05% -23.1% 67.57% 47.75% -29.3%
Llama3.2-3B Base 43.54% 42.34% -2.8% 63.36% 56.76% -10.4%
Empowered 37.84% 43.54% +15.1% 69.67% 57.36% -17.7%
Reflected 43.54% 34.83% -20.0% 63.36% 56.76% -10.4%
SFT 48.05% 37.54% -21.9% 54.65% 48.95% -10.4%
GRPO-MAS-DS-DR 49.25% 43.84% -11.0% 69.97% 63.36% -9.4%
GRPO-MAS-NS-OR 51.95% 45.35% -12.7% 77.78% 61.56% -20.9%
GRPO-nonMAS-DS-DR 49.85% 39.94% -19.9% 77.18% 60.96% -21.0%
GRPO-nonMAS-NS-OR 49.55% 43.54% -12.1% 77.78% 66.37% -14.7%
GRPO-MAS-DS-DR-LConf 44.14% 40.84% -7.5% 75.08% 59.46% -20.8%
GRPO-MAS-DS-DR-LCorr 46.85% 42.34% -9.6% 77.48% 67.27% -13.2%
GRPO-MAS-NS-OR-LConf 48.05% 36.04% -25.0% 77.48% 62.16% -19.8%
GRPO-MAS-NS-OR-LCorr 48.05% 36.04% -25.0% 77.48% 62.16% -19.8%
Llama3.1-8B Base 52.55% 45.35% -13.7% 75.68% 71.47% -5.5%
Empowered 47.75% 43.54% -8.8% 76.88% 72.67% -5.5%
Reflected 52.55% 29.73% $\mathbf{-43.4\%}$ 75.68% 33.03% $\mathbf{-56.3\%}$
SFT 63.36% 50.75% -19.9% 51.05% 50.15% -1.8%
GRPO-MAS-DS-DR 57.96% 51.35% -11.4% 76.88% 76.58% -0.4%
GRPO-MAS-NS-OR 59.46% 59.16% -0.5% 78.38% 72.37% -7.7%
GRPO-nonMAS-DS-DR 52.25% 45.05% -13.8% 75.98% 52.55% -30.8%
GRPO-nonMAS-NS-OR 62.16% 55.56% -10.6% 78.08% 71.17% -8.8%
GRPO-MAS-DS-DR-LConf 55.26% 48.65% -12.0% 76.58% 66.37% -13.3%
GRPO-MAS-DS-DR-LCorr 53.45% 45.65% -14.6% 77.18% 56.46% -26.8%
GRPO-MAS-NS-OR-LConf 54.65% 47.75% -12.6% 77.48% 76.88% -0.8%
GRPO-MAS-NS-OR-LCorr 61.56% 53.15% -13.7% 78.38% 66.07% -15.7%
Llama3.3-70B Base 75.38% 76.28% +1.2% 79.28% 76.58% -3.4%
Empowered 72.67% 76.58% +5.4% 78.68% 78.38% -0.4%
Reflected 75.38% 75.68% +0.4% 79.28% 76.58% -3.4%
Qwen2.5-32B Base 58.56% 60.66% +3.6% 78.98% 69.97% -11.4%
Empowered 56.76% 57.36% +1.1% 79.28% 66.07% -16.7%
Reflected 58.56% 55.56% -5.1% 78.98% 70.87% -10.3%
Qwen2.5-72B Base 58.86% 66.37% +12.8% 78.68% 77.48% -1.5%
Empowered 53.75% 66.97% $\mathbf{+24.6\%}$ 78.98% 78.38% -0.8%
Reflected 58.86% 60.66% +3.1% 78.68% 81.38% +3.4%
GPT-OSS-120B Base 82.88% 82.58% -0.4% 77.48% 73.27% -5.4%
Empowered 81.98% 84.68% +3.3% 80.48% 72.67% -9.7%
Reflected 82.88% 86.19% +4.0% 77.48% 76.88% -0.8%
Gemini-2.5-Pro Base 90.99% 82.58% -9.2% 77.18% 66.07% -14.4%
Empowered 90.69% 93.09% +2.6% 76.28% 73.27% -3.9%
Reflected 90.99% 90.99% +0.0% 77.18% 71.47% -7.4%
GPT-5 Base 96.40% 93.39% -3.1% 81.68% 79.28% -2.9%
Empowered 95.50% 94.59% -1.0% 81.68% 82.88% +1.5%
Reflected 96.40% 94.89% -1.6% 81.68% 83.18% +1.8%
@@
Appendix H LLMs Usage
LLMs were used to polish the writing.